OpenStudy (anonymous):

The length of a rectangle is 6 m, and its width is 14 m. What is the area of half of the rectangle?

OpenStudy (anonymous):

area total = lentgh into breadth

Parth (parthkohli):

The area of half the rectangle is half the area of the rectangle. Does that make sense?

OpenStudy (anonymous):

what is the intire rectangle

Parth (parthkohli):

The area of our rectangle is \(6 \times 14 =84m^2\). Now, the area of half the rectangle is \(0.5 \times 84\).

Parth (parthkohli):

Got it? @tabbykat5

OpenStudy (jiteshmeghwal9):

first find the total area i.e; 6*14=84 now half of the area = 1/2 of 84

OpenStudy (anonymous):

so the answer is 42?

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

thanks for all the help guys :)

Parth (parthkohli):

42 is the answer to the universe.